• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Ibrahim

    Log entry time 08:11:59 on October28,2002

    Entry number 89191


    keyword=Owl Shift Summary

    Owl Shift Summary. New crew on shift: E. Voutier, A. Camsonne, H. Ibrahim.

    12:00 a.m. Continuing run # 2806 from the previous shift.

    05:23 a.m. MCC called to inform us that hall C has requested the halfwave plate be changed to zero and asked us to monitor the beam and see if it is still usable or not!

    06:00 a.m. Beam is still usable.

    06:12 a.m. Finished Q3D_f101 kinematics. Setting the right angle to 37.36 deg. to start the next kineamtics.

    06:41 a.m. Starting Q3D_f201 kinematics by production runs to minimize the movements of the target.

    07:45 a.m. We noticed that espace T5 totals for runs # 2812 and 2813 do not agree with the end of runs T5 totals. It is not clear to us now what the problem is.

    08:00 a.m. End of Owl Shift.

    Continuing Q3D_f101 kinematics:

    Run # 2806 (Production, 1M events)
    Run # 2807 (Production, 1M events)
    Run # 2808 (Production, 1M events)
    Run # 2809 (Production, 1M events)
    Run # 2810 (Production, 1M events)
    Run # 2811 (Production, 1M events)
    Run # 2812 (Production, 1M events)

    Q3D_f201 kinematics:

    Run # 2813 (Production, 1M events)
    Run # 2814 (Production, in progress)